### v1

# Missing Money and Market-Based Adequacy in Deeply Decarbonized Power Systems with Long-Duration Energy Storage

Abstract
The ability of deeply decarbonised power systems to ensure adequacy may increasingly depend on long-duration energy storage (LDES). A central challenge is whether capacity markets (CMs), originally designed around thermal generation, can provide efficient investment signals when storage becomes a central participant. While recent studies have advanced methods for accrediting variable renewables and short-duration storage, the effectiveness of these methods in CMs with substantial LDES penetration remains largely unexplored. To address this gap, we extend a two-stage stochastic equilibrium investment model by endogenising continuous, duration-based capacity accreditation for storage and apply it to a Great Britain-based case using 40 years of weather-driven demand and renewable profiles under varying emission limits. Results show that well-calibrated CMs can sustain near-efficient investment and mitigate revenue volatility, but their effectiveness diminishes in deeply decarbonized systems, underscoring both their potential and the regulatory challenges of supporting large-scale LDES.
Index Terms:
Long-Duration Energy Storage, Missing Money Problem, Capacity Market, Capacity Accreditation
Nomenclature
Sets
$\Omega$
Set of scenarios.
$\mathcal{T}$
Set of time periods.
$\mathcal{G}$
Set of generators.
$\mathcal{S}$
Set of storage units.
$\mathcal{R}$
Set of resources $\mathcal{G}\cup\mathcal{S}$.
$\mathcal{Z}$
Set of durations.
Variables
$c_{g},c_{s}^{P},c_{s}^{E}$
Installed capacities: generator, storage power, storage energy.
$c^{CM}_{g},c^{CM}_{s}$
Capacity market bids: generator, storage.
$q_{\omega tg}$
Generator power generation.
$q_{\omega ts}^{ch},q_{\omega ts}^{dis}$
Storage charging and discharging power.
$d_{\omega t}^{fix},d_{\omega t}^{flex}$
Demand met: price inelastic, flexible.
$e_{\omega ts},e_{s}^{init}$
Storage energy: at time $t$ and initial.
$d^{E}_{\omega,t},d^{C}$
Consumption of energy and capacity .
$\lambda^{E}_{\omega,t},\lambda^{CM},$
Energy, capacity and emission price
$\lambda^{CT}$
Input Parameters
$w_{\omega t}$
Duration of time period [hours].
$\delta_{\omega}$
Probability weight of scenario.
$B$
Benefit factor for meeting demand.
$D_{\omega t}^{fix},D_{\omega t}^{flex}$
Maximum demand: fixed, flexible.
$C_{g}^{V}$
Variable cost of generator.
$\widetilde{I}_{g},I_{g}^{F}$
Generator costs: investment, fixed operational.
$\widetilde{I}_{s}^{P},I_{s}^{P.F}$
Storage power capacity costs: investment, fixed operational.
$\widetilde{I}_{s}^{E},I_{s}^{E.F}$
Storage energy capacity costs: investment, fixed operational.
$\eta_{s}^{ch},\eta_{s}^{dis}$
Storage efficiencies: charging, discharging.
$A_{\omega gt}$
Availability factor of generator.
$EF_{g}$
Emission factor of generator $g$.
$\bar{E}^{CT}$
Annual expected emission limit.
I Introduction
Recent long‐term planning studies have underscored the indispensable role of large‐scale energy storage (ES) in achieving deep decarbonisation of the power sector [1]. Modelling analyses consistently suggest that in futures with very tight greenhouse gas (GHG) emission targets, the additional energy capacity (MWh) requirement is much higher relative to the additional power capacity (MW), indicating that investments in long duration energy storage (LDES) technologies [^1] While LDES does not have a common definition, it usually refers to a class of storage technologies with discharge duration that exceeds 6 hours [2]. will be essential [3]. This need becomes particularly acute in the full or partial absence of alternative clean, firm resources, such as new nuclear power or natural‐gas plants equipped with carbon capture and storage (CCS) [4].
Integration of LDES faces at least two challenges. First, many LDES technologies are capital-intensive and relatively immature. To guide technology development, a recent research stream explored cost thresholds for LDES to appear in least-cost capacity mixes, concluding that significant reductions in energy-capacity costs are needed to compete with other firm capacity resources [5]. Second, the business models that allow LDES to valorize all its operational and environmental benefits for the system [6] have not yet been established. In energy markets with price caps lower than the value of lost load (VOLL), a business model purely based on energy arbitrage remains uneconomic, with narrow price spreads failing to yield sufficient returns to fully support investments [7].
Pure arbitrage is equally unattractive for short-duration energy storage (SDES), e.g., electrochemical batteries. However, in the case of SDES, stacking revenues from multiple market products such as energy and ancillary services [8], combined with generous subsidies, such as tax credits or grants [9], has boosted batteries’ profits and deployment. These two drivers have created robust SDES business models, leading to a sixfold surge of global annual installations since 2020 [10]. However, these models will likely not be transferable to LDES. Short-term applications, such as frequency regulating reserves, and diurnal arbitrage, will continue to be dominated by SDES and additionally face market saturation risks [11]. Furthermore, subsidies and grants can become unsustainable if a large number of clean energy assets needs support.
One benefit of LDES lies in its peaking potential and its contribution to system adequacy [6]. As such, in decarbonized power systems, LDES primarily replaces dispatchable conventional peakers, limits renewable curtailment, and ensures adequacy during contingencies or prolonged periods with low‑VRE generation (so‑called ’energy droughts’ [12]). Consequently, as with peaking units, a substantial portion of the LDES revenues is likely to be realized during scarcity or near-scarcity events. In markets with price caps lower than the VOLL, insufficiently high price spikes during scarcity events could create a missing money problem [13], challenging the viability of investment in LDES. Furthermore, the concentration of scarcity events within particular weather years can amplify interannual profit variability, undermining investor confidence even when wholesale energy price arbitrage provides sufficient revenue over the financial lifetime of the asset [14].
Capacity remuneration mechanisms (CRMs), such as capacity markets (CMs), address missing money and revenue volatility in price‑capped markets. However, CMs were originally designed with thermal generation in mind, and their effectiveness in contexts with substantial ES capacity, particularly LDES, remains understudied. Recent research has focused on accurately valuing the contribution of energy storage to system adequacy. Zachary et al. [15] demonstrated that a change in expected unserved energy (EUE) is a suitable metric for evaluating the contribution of resources to system adequacy, extending the analyses to VRE and storage (yet with strong assumptions about its dispatch). They also demonstrated that a marginal, rather than an average, approach could yield accurate investment signals [^2] Marginal capacity credit reflects the incremental contribution of a new resource to system adequacy, while average capacity credit captures the overall contribution of all existing resources of that type.. Although marginal accreditation methods are more complex to apply, multiple CM operators, including NYISO [16], ISO-NE [17], PJM [18], and GB [19], have recently begun to progress in this direction.
The shift towards marginal accreditation is promising, as investment decisions are highly sensitive to the chosen methodology and underlying assumptions [20]. Yet, even when marginal credits are used for ES and VRE, there is no general proof ensuring that investments will be efficient. Wang et al. [21] demonstrated through simulations that marginal accreditation in CMs can provide price signals that encourage near-efficient investment for VRE and SDES. While these results are encouraging for the effectiveness of CMs in the near future, with narrow participation of ES, they are of limited use for future systems where LDES may play a dominant role.
In general, most existing studies focus on ES with a single value (usually short) of duration or study systems where ES is a minor contributor to a system’s resource adequacy [21, 22]. Many that include ES in CRMs often make simplified assumptions about accreditation [23]. These simplifications make the results of existing studies less relevant for systems with LDES. The duration of LDES, for many emerging technologies, is a design choice independent of the power capacity. Moreover, LDES duration can range from a few to several hundred hours. Different durations yield distinct contributions to system adequacy, complicating decisions for both investors and regulators.
Given the lack of studies evaluating the effectiveness of CMs in providing efficient investment signals for LDES, this article fills this gap by simulating a system with LDES and assessing CM performance, contrasting their results with those of a competitive market and no price cap. We develop an equilibrium model for the electricity market that makes investment and dispatch decisions under a cap on CO2 emissions. We advance current practice by developing and including in our model an accreditation mechanism that explicitly considers capacity credits as a function of ES duration. Our evaluation framework assesses how CMs mitigate the missing‑money problem in terms of both welfare and adequacy. In particular, our study addresses the following questions:
  • How significant is the missing money problem for LDES in price-capped electricity markets under different decarbonisation targets?
  • To what extent can well-calibrated CMs, incorporating marginal accreditation, support efficient investment outcomes?
  • How do inaccuracies in marginal accreditation affect market outcomes and adequacy?
  • To what extent do CMs reduce interannual volatility in energy prices and LDES revenues under different decarbonisation targets?
We contribute to the literature in several ways. Methodologically, we extend the framework of [21] by incorporating a price-elastic demand to address degeneracy and improve price interpretability under high VRE and storage penetration. This is also the first study to endogenise continuous ES accreditation within an investment model, using piecewise linear functions of duration. This improves on previous single-duration methods, which misestimate the adequacy value of storage systems with different durations.
Our analysis offers new insights into market-based adequacy in low-carbon systems with high LDES levels. We confirm that the missing money problem remains significant under price caps, but declines as storage is able to capture wider spreads. We also show that the price cap affects revenues and operations below the cap itself, particularly in cases of low CO2 emissions. Well-calibrated CMs with marginal EUE-based accreditation maintain near-efficient investment signals, but their effectiveness is lower under tighter decarbonization targets.
These findings are relevant for CM design and resource adequacy methodologies. They also inform policymakers considering LDES support mechanisms, such as the UK’s cap-and-floor scheme [24] by highlighting the interplay between accreditation, revenue volatility, and revenue sufficiency.
The rest of this paper is structured as follows. Section II describes the methodology, Section III presents the case study, Section IV discusses the results, and Section V concludes with a synthesis of our findings and suggestions for future research.
II Methodology
This section presents the equilibrium model, including its mathematical formulation, solution strategy, capacity accreditation method, CM parameterisation, and experimental design.
II-A Equilibrium capacity expansion model
Here, we introduce a two-stage stochastic equilibrium model for the investment and dispatch in the electricity market. In the first stage, investments are made, and the capacity market is cleared, and the emission price is decided through a cap-and-trade CO2 scheme. In the second stage, dispatch is decided when the energy market is cleared. To capture variations in load profiles and VRE availability, the second stage simulates dispatch decisions for alternative annual scenarios represented by the set $\Omega$. We model three types of participants: generators, ES operators, and consumers, all of whom are assumed to bid truthfully in a perfectly competitive environment. Although the current setup disregards the network, it can be readily extended to include the system operator.
II-A1 Generators
Let $\mathcal{G}$ be a set of electricity generators. Each generator has three decision variables: its installed capacity ($c_{g}$), the capacity contracted by the CM ($c^{CM}_{g}$), and the electricity produced in each time interval ($q_{\omega,t,g}$), which are in the set $\Gamma^{g}=\left\{q_{\omega,t,g},c_{g},c^{CM}_{g}\geq 0\right\}$. Acting as a price taker, each generator solves the optimization problem (1) below.
$\forall g\in\mathcal{G}\quad\max_{\Gamma^{g}}\sum_{\omega t}w_{\omega t}\delta_{\omega}\left(\lambda^{E}_{\omega,t}-C_{g}^{V}-\lambda^{CT}EF_{g}\right)q_{\omega tg}$
$+c^{CM}_{g}\lambda^{CM}-c_{g}\left(\widetilde{I}_{g}+I_{g}^{F}\right)\hskip 0.0pt$ (1a)
$q_{\omega tg}\leq c_{g}A_{\omega tg}\hskip 0.0pt\quad\forall\omega,t;\quad(\nu^{C}_{\omega tg})$ (1b)
$c^{CM}_{g}\leq CC_{g}c_{g}\hskip 0.0pt$ (1c)
Eq. (1a) is the generators’ objective function, maximising the expected net revenue from participation in the three markets, reduced by the operational and annualised investment costs. Eq. (1b) limits hourly production to the installed capacity multiplied by the hourly capacity factor, while Eq. (1c) limits the CM capacity to the accredited capacity.
II-A2 Storage
Let set $\mathcal{S}$ include all ES market participants. ES participants determine power and energy capacity independently. Additionally, ES decision variables include the capacity contracted by the CM, its initial state-of-charge (SOC), its charging, and discharging. All decision variables are in $\Gamma_{s}=\left\{q_{\omega,t,s}^{\text{dis}},q_{\omega,t,s}^{ch},c_{s}^{P},c_{s}^{E},e_{\omega,t,s},e^{init},c^{CM}_{s}\geq 0\right\}$ and their values are determined by solving the optimization problem (2).
$\forall s\in\mathcal{S}\quad\max_{\Gamma^{s}}\sum_{\omega t}w_{\omega t}\delta_{\omega}\lambda^{E}_{\omega,t}\left(q_{\omega ts}^{\text{dis}}-q_{\omega ts}^{\text{ch}}\right)$
$+c^{CM}_{s}\lambda^{CM}-c_{s}^{P}\left({\widetilde{I}}_{s}^{P}+I_{s}^{P.F}\right)-c_{s}^{E}\left({\widetilde{I}}_{s}^{E}+I_{s}^{E.F}\right)\hskip 0.0pt$ (2a)
$q_{\omega ts}^{\text{dis}}\leq c_{s}^{P}\hskip 0.0pt\quad\forall\;\omega,t$ (2b)
$q_{\omega ts}^{\text{ch}}\leq c_{s}^{P}\hskip 0.0pt\quad\forall\;\omega,t$ (2c)
$e_{\omega ts}=e_{\omega t-1s}+w_{\omega t}(q_{\omega ts}^{\text{ch}}\eta_{s}^{\text{ch}}-\frac{q_{\omega ts}^{\text{dis}}}{\eta_{s}^{\text{dis}}})\quad\forall\;\omega,\;t>1$ (2d)
$e_{\omega ts}=e_{s}^{init}+w_{\omega t}(q_{\omega ts}^{\text{ch}}\eta_{s}^{\text{ch}}-\frac{q_{\omega ts}^{\text{dis}}}{\eta_{s}^{\text{dis}}})\quad\forall\;\omega,\;t=1$ (2e)
$e_{\omega ts}\geq e_{s}^{init}\hskip 0.0pt\;\forall\;\omega,\;t=T$ (2f)
$e_{\omega ts}\leq c_{s}^{E}\hskip 0.0pt\quad\forall\;\omega,t$ (2g)
$c^{CM}_{s}\leq c_{s}^{P}f^{CC}(c_{s}^{E},c_{s}^{P})\hskip 0.0pt$ (2h)
Eq. (2a) is the ES’s objective function, maximising its profit. Eq. (2b)-(2c) limit charging and discharging based on installed power capacity. Eq. (2d)-(2e) calculate the energy stored at the end of a time interval of duration $w_{\omega,t}$ by considering the energy stored at the start of the interval and the charging, and discharging during the interval. Eq. (2f)-(2g) limit the energy stored relative to the energy stored at the start of the model’s horizon and the installed energy capacity, respectively [^3] Constraining end-of-horizon storage to exceed initial levels may be restrictive, but it provides a practical way to handle model end effects under limited foresight. It is also consistent with EU rules requiring minimum gas storage levels before winter [25]..
Finally, (2h) limits the capacity contracted by the CM, $c^{CM}_{s}$, to the qualifying capacity of ES. We determine the capacity credit of ES in a novel manner as a function, $f^{CC}$, of ES duration $\zeta_{s}$ $\eta_{s}^{\text{dis}}c_{s}^{E}/c_{s}^{P}$. This new way allows us to properly consider the capacity market contributions of a range of ES durations. The function $f^{CC}$ is piece-wise linear with $|L|$ segments, where each segment $l$ covers a range of durations $\zeta_{s,l-1},\zeta_{s,l}]$. As such, $f^{CC}$ is represented as
$f^{CC}(\zeta_{s})=\alpha_{ls}+\beta_{ls}\zeta_{s},\quad\forall\;\zeta_{s}\in[\zeta_{sl-1},\zeta_{sl}],l\in\mathcal{L}$ (3)
Capacity credit curves are monotonically increasing and concave with respect to duration [26]. In other words, the marginal contribution to adequacy diminishes as duration increases (i.e., the value of $\beta_{l,s}$ is positive but lower for higher durations). We estimate the coefficients of $f^{CC}$, i.e., $\alpha_{l,s},\beta_{l,s}$, using piece-wise linear regression.
In the maximisation problem (2), constraint (2h) can be included as is because $f^{CC}$ is concave. If $f^{CC}$ was not concave, binary variables and associated constraints would be needed to ensure that the segments of the piecewise curve are activated in a logical order. Multiplying both sides of (3) with the installed power capacity $c_{s}^{P}$, (2h) becomes:
$c^{CM}_{s}\leq\alpha_{l,s}c_{s}^{P}+\beta_{l,s}\eta_{s}^{\text{dis}}c_{s}^{E},\quad\forall l\in\mathcal{L}$ (4)
II-A3 Consumers
We assume a single representative agent acting on behalf of all consumers. Its decision variables include the electricity consumption and capacity purchased in the CM, grouped in the set $\Gamma^{C}=\left\{\left.d^{E}_{\omega,t}\right|t\in\mathcal{T},\omega\in\Omega,d^{C}\geq 0\right\}$. Consumers maximise the expected consumer surplus, which is defined as the difference between the integral of willingness to pay (WTP) functions for energy $f^{E}$ and capacity $f^{C}$ and the procurement cost (product of price and quantity purchased).
$\max_{\Gamma^{C}}\sum_{\omega,t}w_{\omega t}\delta_{\omega}\Bigg{[}\int_{0}^{d^{E}_{\omega,t}}f^{E}_{\omega,t}(x)dx-\lambda^{E}_{\omega,t}d^{E}_{\omega,t}\Bigg{]}\\ +\int_{0}^{d^{C}}f^{C}(x)dx-\lambda^{C}d^{C}\hskip 0.0pt$ (5a)
The demand curves for electricity can take diverse shapes, ranging from perfectly price-inelastic to perfectly price-elastic curves. In this work, in line with [27], we assume that the demand curve for electricity at each time interval has two segments. The first segment is perfectly elastic and demands electricity between 0 and $D^{fix}_{\omega,t}$ with a willingness to pay up to $B$. The second segment is a price-elastic linear demand curve where the willingness to pay ranges from 0 to $B$ for consumption between 0 and $D^{flex}_{\omega,t}$. This demand curve helps ensure price uniqueness and resolve dual degeneracy issues in systems with high shares of renewables and storage. The WTP function for energy $f^{E}$ is included in the model using the equations below.
$\int_{0}^{d^{E}_{\omega t}}f^{E}_{\omega t}(x)dx=B\left(d_{\omega t}^{\text{fix}}+d_{\omega t}^{\text{flex}}-\frac{(d_{\omega t}^{\text{flex}})^{2}}{2D_{\omega t}^{\text{flex}}}\right)\hskip 0.0pt$ (6a)
$d_{\omega t}^{\text{fix}}\leq D_{\omega t}^{\text{fix}}\quad\forall\omega t\hskip 0.0pt$ (6b)
$d_{\omega t}^{\text{flex}}\leq D_{\omega t}^{\text{flex}}\quad\forall\omega t\hskip 0.0pt$ (6c)
The CM function is represented similarly, and it often has more segments with differing levels of elasticity. The CM function $f^{C}(d^{C})$ represents the WTP for capacity, and is a monotonically decreasing function. It consists of $N$ segments, with differing levels of elasticity. The first segment is perfectly price-elastic, willing to pay the capacity market price cap $B^{C}_{1}$ for any capacity between 0 and $D^{\text{C-fix}}$. In the last segment, the WTP $B^{C}_{N}$ decreases to zero as $d^{C}$ reaches the maximum capacity the CM aims to procure. The N2N-2 segments in between are flexible (lengths $D_{n}^{\text{C-flex}}$, starting prices $B^{C}_{n}$, with $B^{C}_{N}=0$) and they are parametrized based on net-CONE. Given this structure, the function can be represented as:
$\int_{0}^{d^{C}=d^{\text{fix}}+\sum_{n=1}^{N-1}d_{n}^{\text{flex}}}f^{C}(x)\,dx=B^{C}_{1}d^{\text{C-fix}}+\\ \sum_{n=1}^{N-1}B^{C}_{n}d_{n}^{\text{C-flex}}-\sum_{n=1}^{N-1}\frac{(d_{n}^{\text{C-flex}})^{2}(B^{C}_{n}-B^{C}_{n+1})}{2D_{n}^{\text{C-flex}}},\hskip 0.0pt$ (7)
where $d^{\text{C-fix}}$ is the level procured for the segment that is willing to pay up to the capacity market price cap, and $d_{n}^{\text{C-flex}}$ is the capacity procured for the $n$-th flexible segment. Our model includes segment-specific constraints similar to (6b) - (6c).
II-A4 Market clearing conditions
Finally, agent-specific decisions must comply with system-wide constraints. In particular, market-clearing conditions ensure a balance between supply and demand for both energy and CMs, and a constraint on the annual expected emissions ensures that the CO2 emissions from electricity generation do not exceed the cap. The dual variables of these constraints, $\lambda^{E}_{\omega,t}$, $\lambda^{C}$, and $\lambda^{CT}$, can be interpreted as market prices (weighted with scenario probabilities and time interval durations).
$\sum_{g}q_{\omega t,g}+\sum_{s}\left(q_{\omega ts}^{\text{dis}}-q_{\omega ts}^{\text{ch}}\right)$
$=d_{\omega t}^{\text{fix}}-d_{\omega t}^{\text{flex}}\quad\forall t,\omega;\hskip 0.0pt(\lambda^{E}_{t,\omega})$ (8a)
$\sum_{g}c^{CM}_{g}+\sum_{s}c^{CM}_{s}=d^{CM};\hskip 0.0pt(\lambda^{CM})$ (8b)
$\sum_{g,t,\omega}w_{\omega t}\delta_{\omega}q_{\omega t,g}EF_{g}\leq\bar{E}^{CT};\hskip 0.0pt(\lambda^{CT})$ (8c)
II-B Solution strategy
The market players’ problems are convex quadratic with linear constraints, making their Karush–Kuhn–Tucker (KKT) conditions necessary and sufficient for optimality [28]. By combining KKT conditions and applying the integrability principle of variational inequalities [29], the solution to the central planner’s convex quadratic optimization problem equates to the equilibrium solution [29].
II-C Capacity accreditation
We estimate capacity credits using the marginal accreditation approach, supported by academic literature and increasingly adopted in industry [15]. Credits are calculated for each generator $g\in\mathcal{G}$ and storage unit $s\in\mathcal{S}$, with storage evaluated across various durations $\zeta\in\mathcal{Z}$. Thus, the resource set for accreditation is $\mathcal{R}=\mathcal{G}\cup(\mathcal{S}\times\mathcal{Z})$.
The accreditation methodology follows similar principles to those proposed by the NESO for marginal EFC [19] and builds upon the structure presented by Wang et al. [21] to consider multiple storage durations. We first determine the baseline EUE resulting from dispatching the welfare-maximising capacity mix under the price-capped market. Unlike in conventional models with perfectly inelastic demand, the definition of unserved energy is less straightforward when demand is price-elastic [^4] It is also worth noting that the solution to our equilibrium problem, while maximising social welfare, may not be EUE minimising. EUE-minimising storage dispatch is one of the assumptions in Zachary et al. [15] for integrating ES into capacity markets, which may not hold in our framework, as the storage dispatch is driven by economics.. In this work, we follow an approach similar to [30]. Considering all demand with WTP greater than or equal to the price cap as inelastic, we record any shortfall of consumption with respect to that demand as unserved [^5] Note that the PC not only limits the energy price levels during scarcity but also removes part of the flexible demand, which would voluntarily adjust its level based on the energy price.. The unserved energy $l_{\omega t}$ is given by:
$l_{\omega t}=\max(0,D_{\omega t}^{\text{fix}}+\frac{D^{\text{flex}}_{\omega t}}{\text{VOLL}}\cdot(\text{VOLL}-\text{PC})-d_{\omega t}^{\text{fix}}-d_{\omega t}^{\text{flex}})$ (9)
Then, the EUE is calculated as:
$\text{EUE}=\sum_{\omega t}w_{\omega t}\delta_{\omega}l_{\omega,t}$ (10)
Subsequently, we perturb the capacity mix by adding a hypothetical perfect generator with small capacity $\epsilon$ and re-solve the model to obtain the reference EUE. Then, for each resource, including storage with specific durations, we add $\epsilon$ capacity of this resource to the mix and calculate the capacity credit as the ratio of the EUE reduction to that of the perfect generator. The complete process is presented in Algorithm 1.
Algorithm 1 Capacity Accreditation
  Input: Optimal capacity mix $\hat{c}_{r},\forall r\in R=G\cup(S\times E)$; system data; small capacity addition $\epsilon>0$; duration set $\mathcal{Z}=\{\zeta_{1},\zeta_{2},\ldots\}$; model $\mathcal{M}$ composed of (1),(6),(8), with a price-capped energy demand function.
  Output: Capacity credit $CC_{r},\forall r\in R$
  Step 1: Compute baseline EUE
  Fix $c_{r}\leftarrow\hat{c}_{r},\forall r\in R$
  Solve model $\mathcal{M}$ to obtain baseline $\text{EUE}_{0}$
  Step 2: Perturb capacity for perfect generator
  Add $\epsilon$ of a perfectly flexible, reliable reference generator: $c_{\text{ref}}\leftarrow\epsilon$
  Solve model $\mathcal{M}$ to obtain $\text{EUE}_{\text{ref}}$
  Step 3: Perturb capacity for each resource
  for $r\in R$ do
   if $r\in S\times\mathcal{Z}$ then Fix duration of storage $r$ to $\zeta\in\mathcal{Z}$
   Add $\epsilon$ of resource $r$: $c_{r}\leftarrow\epsilon$
   Solve model $\mathcal{M}$ to obtain $\text{EUE}_{r}$
   Compute capacity credit: $CC_{r}=\frac{\text{EUE}_{0}-\text{EUE}_{r}}{\text{EUE}_{0}-\text{EUE}_{\text{ref}}}$
  end for
II-D Capacity market parameters
We construct the capacity demand curve by determining the capacity target and the corresponding WTP function. To determine the WTP, we rely on the system-wide net CONE and an assumed elasticity of demand for capacity. Following [31], we calculate the capacity target $CT$ using the target technology-specific capacity installations and the estimated capacity credits.
$CT=\sum_{g}CC_{g}\dot{c}_{g}+\sum_{s}f^{CC}(\dot{\zeta}_{s})\dot{c}_{s}^{P}$ (11)
Net CONE is defined as the difference between a technology’s annualised fixed and operational costs and its expected net revenues from the energy market, including ancillary services where applicable. Traditionally, CMs estimate net CONE for a reference technology, typically a flexible conventional generator whose accredited capacity is based on its availability. We follow industry practice and use CCGT as the reference technology, assuming, for demonstration purposes, its perfect availability and flexibility.
Assuming that the price cap distorts prices only during hours with scarce supply and that the CC captures the impact of the price cap distortion during those hours on the financial bottom line of a resource, the net CONE per accredited MW in a long-term equilibrium should theoretically be the same for all resources [32]. Anticipating this long-run outcome, we estimate alternative values for CC, hereafter referred to as required CC, with a clear economic interpretation [33]. Eq. (12) computes the required $CC^{\text{req}}_{r}$ as the ratio of the net CONE per installed MW of technology $r$ ($\text{NC}^{\text{uc}}_{r}$) divided by the net CONE per installed MW of the reference technology ($\text{NC}_{\text{ref}})$666 The required CC is valid only for resources that have positive installed capacity in the social-welfare maximizing capacity mix. Only these resources should recover their costs from the capacity and energy markets in the long run.. Eq. (12) assumes that the accredited capacity of the reference technology is equal to its installed capacity.
$\text{NC}_{\text{ref}}=\frac{\text{NC}^{\text{uc}}_{r}}{CC^{\text{req}}_{r}}\rightarrow CC_{r}^{\text{req}}=\frac{\text{NC}^{\text{uc}}_{r}}{\text{NC}_{\text{ref}}}\quad\{\forall r\in\mathcal{R}|\hat{c}_{r}>0\}$ (12)
II-E Experiment design
For our simulations, we use the model presented in II-A in two distinct market structures: an energy-only market (EOM), where CM-related variables and constraints are omitted, and a combined energy and CM (E+CM). Each structure can be run with different energy demand functions, allowing the price to spike to VOLL or be capped at the PC, set at an administratively defined level significantly lower than VOLL. When prices are allowed to spike to VOLL, the amount of the price-elastic demand is $D^{\text{flex}}_{\omega,t}$. With the PC, the width of the linear part of the demand curve reduces by $D^{\text{flex}}_{\omega,t}\cdot\frac{\text{VOLL}-\text{PC}}{\text{VOLL}}$ and the flat part of the demand curve increases in width by the same amount.
Effectively, we run the following main combinations: EOM VOLL is an energy-only market WTP function rising to VOLL; EOM PC is an energy-only market with WTP capped at PC; E+CM is an energy+capacity market with WTP capped at PC. For comparative and accreditation purposes, we also run a case denoted EOM PC opt. mix, where the price is capped, but the capacity mix is fixed to the welfare-maximising mix from EOM VOLL.
The welfare-maximising capacity mix is determined by solving the EOM VOLL model. Next, the EOM PC opt. mix model provides us with the baseline level for $\text{EUE}_{o}$, which is used to calculate capacity credits. Note that when the price cap is applied, the welfare-maximizing mix is no longer economically feasible due to limited price spikes during scarcity, which curb the cost recovery. Finally, to identify financially viable capacity mixes in the presence of a cap on energy prices, we solve the E+CM and EOM PC models, which include and lack a capacity market, respectively.
II-F Performance metrics
We use two primary performance metrics to evaluate the effectiveness of the E+CM model. The first metric is the social welfare, which is calculated as the difference between the consumer benefit (6a) and capital and operating costs. By construction, the SW of the EOM VOLL run will be the highest because in that case, the objective function of the optimisation problem is identical to the SW metric. The SW lost in the other runs measures the cost of a price cap or the effectiveness of the E+CM design in terms of SW. The second metric is EUE, as defined earlier in (10). In addition to our main metrics, we discuss how different the E+CM capacity mix is compared to the welfare-maximising capacity mix.
In this methodological framework, three factors can contribute to social welfare inefficiencies in the E+CM market. First, calibrating CM parameters is challenging, and estimation errors in capacity credits or net-CONEs can impact the results. Second, imposing price caps narrows price spreads, disrupting storage operational patterns. Third, the unserved demand may be inefficiently allocated between periods as the model assigns a constant value of PC to every MWh unserved. We evaluate SW according to the true WTP function, following Joskow and Tirole [34] in assuming efficient (non-price) rationing so that customers with the highest WTP are prioritised, but this means the average per-MWh cost of unserved energy rises with the amount unserved in that period. Under the alternative assumption of random rationing, the welfare cost of every MWh unserved would be equal, but the total cost of any given reduction in load would be higher than if it was targeted at the lower-WTP customers.
The latter two effects above relate solely to the introduction of a PC, and can be assessed by running the EOM PC case with optimal mix obtained from EOM VOLL (EOM PC opt. mix. mentioned earlier). To isolate these effects, we need to obtain the optimal true WTP-aware discharge schedule for the storage and remove the effects of true WTP inobservability on the EUE distribution, i.e., isolating the third effect. First, we fix the generation and charging decisions resulting from the original run. Subsequently, we replace the price-capped WTP with the true WTP, allowing it to reach the VOLL. Finally, we redispatch the mix, with only the discharging variables unfixed, which will now be scheduled to maximise true SW. Because the charging decisions of the storage are fixed, the schedule will only change discharging to better allocate the EUE across the time steps. Through this adjustment, we are left with only inefficiencies related to the first two effects, which are relevant to investigation in this paper. The third is out of scope as it relates to how to incentivize effective integration of demand-side flexibility under a price cap. Hence, in this paper, we assume that another scheme has successfully integrated demand-side flexibility to effectively distribute unserved energy.
III Case study setup
[…]
We study a diverse set of near-zero emission systems, which we identify by varying annual EI limits from 15 to 0.1 gCO2/kWh [^7] The selected EIs align with recent Future Energy Scenarios [39]. The model includes emission caps, which are calculated as the product of EI and the expected annual demand ($\sum_{t,\omega}[w_{\omega,t}\delta_{\omega}(D_{\omega,t}^{fix}+D_{\omega,t}^{flex})]$). Assuming 90% CCS efficiency, CCGT with CCS emits 37.8 gCO2/kWh. LDES is excluded from the optimal mix without an emission cap, concurring with previous research on the need for technological advancements [5].
We construct the demand curves for the CM as follows. Following GB CM guidance [40], the curve has two downward-slopping segments, and their breakpoints are: (0.965•$CT$, 1.5•net-CONE), ($CT$, net-CONE), (1.035•$CT$, 0). The target capacity $CT$ is established as defined in II-C.
TABLE I: Techno-economic Parameters of Considered Technologies
Parameter CCGT CCS Solar Wind Battery LDES
Power Capex (\$/kW) 2500 895 1335 306 2000
Energy Capex (\$/kWh) - - - 223 10
FOM (\$/kW-yr) 27 15 28 7.6 (5.6) 40 (0.1)
Fuel cost + VOM (\$/MWh) 40 0.5 0.5 0.5 0.5
Charge Eff. (%) - - - 92 60
Discharge Eff. (%) - - - 92 50
Lifetime (years) 30 30 30 15 30
WACC (%) 7.1 6.2 6.2 7.1 7.1
FOM costs per MWh of storage energy capacity.
IV Results
This section presents the results of the CM analysis, evaluating the E+CM design against the EOM benchmark across varying EI scenarios. Organised into five subsections, it examines the missing money problem, capacity credit estimation, the CM effectiveness and its effects on LDES, and sensitivity to accreditation inaccuracies.
IV-A The missing money assessment
We derive a benchmark mix from EOM VOLL simulations for each EI case. Fig. 1 shows the capacity mix and LDES duration. Tighter EI targets reduce CCGT-CCS capacity while increasing wind, solar, and LDES. Batteries play a minor, auxiliary role [^8] We do not model ancillary services, where batteries are expected to be essential.. Despite technological changes, non-VRE capacity remains near 80% of peak demand, as firm capacity needs slowly decline with VRE expansion. Fig. 1b concurs with [3], with LDES duration rising sharply as decarbonisation deepens.
Figure 1: Capacity results for the ‘EOM with VOLL’ simulations. a) shows the installed power capacity as a percentage of peak demand, while b) shows the installed duration of LDES.
Fixing capacities at the levels from Fig. 1, we dispatch resources under a PC to estimate how much money each technology is missing to break even. Fig. 2 shows net revenues from the energy market per technology for two sets of time steps that have energy prices Below and Above or at PC, respectively. For storage, we estimate the cost of energy discharged based on the average charging cost of stored energy at the time of discharge. Appendix 1 describes its calculation.
CCGT-CCS recovers more than 80% of its costs during intervals with prices above the PC in all EI cases. In contrast, storage recovers between 30–80% in these intervals, with this fraction decreasing as emission limits tighten. This trend translates into a diminishing missing money problem for storage. Lower charging prices, driven by higher VRE generation and curtailment, result in reduced storage costs. At the same time, rising carbon costs increase the SRMC of CCGT-CCS, raising discharging prices when CCGT-CCS is the marginal producer, and thereby widening arbitrage spreads. We also observe that, unlike systems without storage, the introduction of the PC not only reduces the Above or at PC net revenues, but also affects net revenues during hours when prices are below the PC, highlighting the intertemporal nature of the missing money problem recently emphasized by [41]. VRE technologies recover a large share of their costs in all cases and are less exposed to missing money. In contrast, CCGT-CCS remains the most vulnerable, as caps suppress scarcity rents that would otherwise support cost recovery.
Figure 2: Cost recovery per technology for EOM VOLL and EOM PC opt. mix.
IV-B Capacity credit estimates
We estimate the marginal CCs presented in Fig. 1 following Algorithm 1. For storage, we compute up to 56 marginal CCs, [^9] For less stringent EI limits, fewer point estimates are required, as a CC of 1 is reached at shorter durations. with one CC per ES duration. We fit a piecewise linear curve with four segments to the CC estimates. The curve closely follows the estimates, yielding an $R^{2}$ value exceeding 0.99 in all cases, which confirms the appropriateness of the piecewise linear approach. Fig. 3 illustrates a critical dynamic: marginal CC estimates are highly sensitive to the underlying resource mix. In general, storage requires longer durations to achieve full accreditation as EI limits become more stringent and dispatchable capacity diminishes. The marginal CC for batteries is higher than for LDES for equivalent durations, owing to better operational characteristics, primarily higher charging efficiency, facilitating greater utilisation under narrower charging/discharging price spreads. We observe that the curves for the 1 and 0.1gCO2/kWh cases are surprisingly close and even cross in the Battery case. To interpret and relate the results on CC with the missing money problem, we contrast the estimated CC to the required CC for all technologies in Fig. 4a, where we also present the net CONEs per marginally accredited MW.
Figure 3: Marginal capacity credit estimates as a function of storage duration. Markers indicate credit obtained with duration from EOM VOLL.
For all cases except 1gCO2/kWh, we observe that the estimated credits are close to the required values and that net CONEs are similar among technologies. To explain the differences in the 1gCO2/kWh case, we examine how ES influences the estimated CC under three paradigms for storage dispatch in Step 2 of algorithm 1. In the first paradigm, labelled unconstrained in Fig. 4b, the dispatch (model M) decides storage charging and discharging. This is the assumption that we have used so far in the article to estimate CC. In the second paradigm, labelled existing ES charging fixed in Fig. 4b, the dispatch decides storage discharging and keeps the ES charging at levels identified by the SW-maximising problem. Last, in the third paradigm, labelled existing ES charging and discharging fixed in Fig. 4b, the dispatch keeps both charging and discharging variables of ES at the levels identified by the SW-maximising solution.
Figure 4: a) Credited net-CONEs and capacity credits for technologies identified by ‘EOM with VOLL’ simulations. Note for storage, the optimal duration is considered. b) CCGT-CCS marginal 0.01MW unit contribution to EUE reduction under alternative setups.
The difference in EUE reduction (output of Step 2 in Algorithm 1) between the second and the third paradigm is expected. Zachary et al. [15] have shown that when a marginal amount of a generator is added to the mix $\hat{c}_{r}$, the reduction in EUE is greater when the mix already contains storage. This is because once the generator addresses issues in hours that storage was previously discharging, storage discharging can be shifted to other hours with a shortfall of energy.
However, the difference in EUE reduction between the first and second paradigms has not been widely discussed before. That difference is more pronounced in the 1gCO2/kWh case. In that case, the CCGT-CCS capacity is so low that it sometimes generates at full capacity to charge storage; yet even then, storage does not always reach the energy level required to address energy shortfalls in time. Hence, when marginal CCGT-CCS is added, part of the EUE reduction should be attributed to storage discharging due to higher levels of stored energy. This dynamic explains why the estimated credit for storage is lower than its required credit. The same phenomenon explains why VRE’s estimated credit is higher than its required credit, as it partially captures storage’s contributions.
The differences between the first two paradigms are negligible in other cases. In the 15 and 5 gCO2/kWh cases, the CCGT-CCS capacity is relatively high compared to ES, and as such, it is not generating at full capacity when it charges storage. In the 0.1gCO2/kWh case, the emission cap is so low that the effective marginal cost of the CCGT-CCS is greater than $\text{PC}\cdot\eta_{s}^{\text{ch}}\cdot\eta_{s}^{\text{dis}}$.
IV-C Effectiveness of energy plus capacity market
Under a price cap, SW losses are caused by two factors. First, unserved demand is inefficiently distributed among consumers due to a lack of price signals between the PC and VOLL. Second, distorted price signals result in inefficient ES dispatch. In Table II, we show SW losses caused by the second factor for the EOM PC and E+CM. For the benchmark capacity mix dispatched under price cap, we present SW losses for two cases: when both factors are present (EOM PC opt. mix. ineff. dist.) and when only the second factor affects results (EOM PC opt. mix.). Under higher emission limits, welfare losses almost disappear when demand rationing is perfect, indicating that inefficient distribution of unserved energy is the primary reason for SW losses. However, for the 0.1gCO2/kWh, the welfare loss is high even under perfect demand rationing, indicating a substantial impact of PC on storage operations via reduced price spreads.
TABLE II: Social Welfare and Underved Energy for All Runs
EI USE [% of total demand] Welfare Loss [% TC in EOM VOLL]
EOM
VOLL
EOM PC
opt. mix
E+CM
EOM
PC
EOM PC
opt. mix.
ineff. dist.
EOM PC
opt. mix
E+CM
EOM
PC
15.0 0.0043 0.0043 0.0044 0.0143 -0.0140 -0.0003 -0.0055 -0.8536
5.0 0.0034 0.0034 0.0036 0.0133 -0.0521 -0.0078 -0.0158 -0.7101
1.0 0.0018 0.0019 0.0029 0.0088 -0.0338 -0.0114 -0.0675 -0.4253
0.1 0.0027 0.0080 0.0080 0.0294 -0.7876 -0.6862 -0.3773 -1.6923
In order to make the results comparable, for the EOM VOLL case we measure EUE with relation to the fixed part of the true WTP and the ’cut off’ part of the flexible demand above the PC.
Under the E+CM design, we observe welfare losses, but they are significantly lower than those observed with PC but no CRM (EOM PC). For all cases with EI higher than 0.1gCO2/kWh, welfare losses are minimal (below 0.1% of total cost) under the E+CM design, yet growing with a tighter emission limit. However, in the 0.1gCO2/kWh case, the welfare loss is almost 0.4% of total cost under the E+CM design, which might seem high, but it is still roughly half the loss caused by the PC alone (EOM PC opt. mix.). In E+CM design, players account for PC distortions, and their optimal solutions differ from the benchmark mix, primarily in terms of reduced LDES capacity (1.3 MW/66 MWh). That mix is more efficiently dispatched, and the SW is closer to the optimal level under the E+CM compared to a design that supports the benchmark capacity mix. At 1gCO2/kWh, losses are primarily due to misaligned credits, which result in a misaligned mix. In terms of EUE, comparing EOM VOLL and EOM PC opt. mix., we see that the impact of the PC is substantial under the lowest EUE scenario. EUE levels under E+CM closely match the EOM PC opt. mix. levels, except for the 1gCO2/kWh case, which again derives from misaligned credits.
IV-D CM impacts on prices and LDES net revenues
Introduction of CM affects energy prices and LDES revenues. Table III contains consumer cost and price volatility metrics. Tighter emission limits raise energy costs. Due to price caps and payments through CMs, energy prices in E+CM are lower by 15–30 \$/MWh compared to EOM, but capacity procurement costs, passed to consumers, offset these savings. Thus, consumer costs are similar in EOM and E+CM, indicating no extra financial burden for the CM. [^10] Administrative and transaction costs, which could marginally impact overall costs, are not considered in this analysis. CMs also lower energy price volatility, though this effect diminishes with stricter emission limits.
TABLE III: Price and Cost Metrics for All Runs
Em. Intensity [gCO2/kWh]
EOM
E+CM
$\mu_{\text{Price}}$ $\sigma_{\text{CV}}$ $\beta_{\text{EB}}$ $\mu_{\text{Price}}$ $\sigma_{\text{CV}}$ $\kappa_{\text{CCC}}$ $\beta_{\text{EB}}$
15.0 100.480 0.410 15.932 80.564 0.240 19.075 15.944
5.0 122.880 0.516 18.171 106.145 0.334 16.128 18.238
1.0 139.682 0.772 11.626 130.229 0.638 9.144 11.707
0.1 157.523 1.249 7.698 134.804 1.019 21.363 6.976
$\mu_{\text{Price}}$: Annual average energy price [\$/MWh]. $\sigma_{\text{CV}}$: Coefficient of variation of annual energy price [–]. $\kappa_{\text{CCC}}$: Average annual consumer capacity cost [\$/MWh]. $\beta_{\text{EB}}$: Emission benefit from carbon credits [\$/MWh].
Note: The emission benefit ($\beta_{\text{EB}}$) derives from the shadow price of the emission constraint. It may not equate to a societal valuation of avoided emissions.
LDES breaks even on average but not annually, as value accrues during rare scarcity events. Fig. 5 shows the annual net revenue distributions for LDES. In all cases, the distribution is highly positively skewed, with more than half of the outcomes reflecting annual losses, but a few extreme cases of very high profits. For cases with tighter emission limits, the CM proves highly effective in reducing both net revenue volatility and the positive skewness, which is an essential consideration for risk-averse investors. However, this stabilising effect weakens as emission constraints tighten.
Figure 5: Distribution of annual net revenues for LDES across EI cases.
IV-E Sensitivity to capacity credit estimation inaccuracies
We assess the impact of capacity credit estimation errors on market performance through a sensitivity analysis, adjusting the capacity credit of storage technologies to account for the dependence of CC on the resource mix and potential inaccuracies of the accreditation methodologies. For each simulation, we use the adjusted credits to estimate a new capacity target. We keep the net-CONE for the reference technology the same across all runs. Fig. 6 shows the change in the average energy price, capacity price, social welfare, and EUE. Under credit underestimation, the investment signal provided by the capacity market to storage weakens, leading to lower investment and tighter conditions in the energy market. As a result of lower capacity targets, capacity market prices reduce, and energy prices and EUE increase. The changes are opposite under credit overestimation. We also find an asymmetric effect on welfare, with underestimation causing larger losses. The SW marginal increases with overestimated credits for 1 and 0.1 gCO2/kWh cases. In the former case, this is due to misaligned credit (per Section IV-B), and in the latter, due to the mix that better negates the impact of the price cap on storage dispatch.
Figure 6: Sensitivity of performance to storage credits.
V Discussion and conclusions
This study investigates whether CMs can effectively address the problem of missing money, caused by price caps, and ensure system adequacy and high social welfare in deeply decarbonised electricity systems with high LDES penetration. Methodologically, we extend state-of-the-art equilibrium investment models by introducing continuous, duration-based accreditation for storage. By formulating capacity credit curves as piecewise linear functions of storage duration, we enable efficient integration of flexible-duration ES into an investment model. Formulations presented here are useful not only for market-based studies but also for reliability-constrained capacity expansion models with planning reserve margins.
The analysis is performed for a case study with renewable and load profiles that resemble those of Great Britain. Simulations consider a range of CO2 emission targets that yield varying levels of LDES integration. We find that price cap distortions result in a substantial missing money problem for LDES (approximately 15% to 50% of LDES costs are not recovered when price caps are retroactively applied to an optimal capacity mix). The magnitude of the problem reduces under more decarbonized systems as storage can capture higher spreads due to low charging costs during periods of abundant renewable supply and high discharging costs during periods with expensive carbon-emitting marginal suppliers.
We demonstrate that well-calibrated CMs can provide near-efficient investment signals, resulting in welfare outcomes comparable to those of an energy-only market without price caps. These results expand the work of Wang et al. [21], by showing that efficient outcomes are possible even when demand is flexible, storage duration is an investment decision, and ES makes up a significant share of installed capacity. In cases of deeper decarbonization, we find that the capacity market is less effective in restoring system security of supply, as measured by expected unserved energy, to the levels that would prevail without price caps. Two factors explain this result. First, the capacity credits required for cost recovery are not always aligned with those estimated through a marginal accreditation approach, as the latter is highly sensitive to assumptions about storage dispatch. Second, in the presence of storage, energy price caps distort market outcomes beyond periods of scarcity. By capping discharging revenues, the price cap undermines incentives for charging in non-scarcity hours. This suggests that conventional methods that only consider the difference between PC and VOLL may be poorly suited to parameterize capacity markets in deeply decarbonized systems.
Our work provides a series of timely regulatory insights. While CMs can, in principle, remain an effective tool for ensuring adequacy, they are complex and sensitive to parameter values such as capacity credits. Nonetheless, CMs effectiveness can remain high in the medium term, provided that accreditation methods are transparent and predictable, giving investors greater certainty. In the longer term, the definition of missing money itself may need to be reevaluated, as it should account for how price caps distort storage revenues across both scarcity and non-scarcity periods. Given that the ability of CMs to reduce price and revenue volatility declines with deeper decarbonisation, complementary measures, such as risk-hedging government-backed contracts, are worth further investigation.
Future work on business models for LDES can extend the framework presented in this article by modeling risk-averse investors, considering relevant policy mechanisms and novel electricity market products that might be a good fit for LDES, and limiting the horizon within which LDES has foresight into future energy prices.
Appendix 1
To estimate storage net revenues per time step, we first calculate the time-varying book value of stored energy based on charging prices. For each scenario $\omega$ and time step $t$, the book value per MWh of stored energy, $\phi_{\omega,t,s}$, is updated as follows:
$\phi_{\omega ts}=\;$ $\left(\phi_{\omega,t-1,s}\cdot\left(e_{\omega,t-1,s}-w_{\omega t}\cdot q^{\text{dis}}_{\omega ts}/\eta_{s}^{\text{dis}}\right)\right$.
$\left.+\;\lambda^{E}_{\omega t}\cdot w_{\omega t}\cdot q^{ch}_{\omega ts}\right)\cdot\left(1/e_{\omega ts}\right)\quad\forall\omega ts\hskip 0.0pt$ (13)
We need to set the value of the initial stored energy. Since we constrain the initial and final energy levels to be equal, as if the model were continually cycling through its year of time periods, we impose a similar restriction on their values. We iterate the calculation of equation (13) until $\phi_{\omega,0,s}=\phi_{\omega,T,s}$, starting from $\phi_{\omega,0,s}=0.$
Considering this time-varying value as the cost of discharging, we estimate the net revenue from participation in storage in the energy market at each time step as follows:
$\pi^{\text{op}}_{\omega ts}=q^{\text{dis}}_{\omega ts}\cdot\left(\lambda^{E}_{\omega t}-\phi_{\omega,t-1,s}/\eta_{s}^{\text{dis}}\right)\quad\forall\omega ts\hskip 0.0pt$ (14)